# A Meat-Summer Night’s Dream: A Tangible Design Fiction Exploration of Eating Biohybrid Flying Robots

### v1

@@ 4.2. Tangible Multisensory Dining Experience as a Valuable Provocation
Our multisensory method surfaced a rich range of perspectives and tensions. Participants connected technological speculation with reflections on politics, religion, nationality, popular culture, symbolism, and ontological beliefs about life, consciousness, and suffering. Speculative dining thus functioned as more than entertainment: it operated as an epistemic tool for engaging publics in complex biotechnological debates. In the reflections that follow, we identify three mechanisms through which these deep speculative exchanges emerged.
4.2.1. The Value of Sensory Immersion and Emotive Response
@@
Participants felt that the immersive dining format—particularly the use of napkin hoods and live/mechanical elements—provoked strong sensory engagement and elicited intense emotional responses. Several highlighted the hooded moment as especially impactful, though their interpretations diverged. P2 described how covering heads with napkins transformed the room: conversation ceased and a collective silence emerged, “like putting a cover over a birdcage.” Rather than diminishing sociality, this paradoxically heightened his awareness of others: he became “hyper-aware of the other people eating around me,” sensing a subtle connection through shared quietude. By contrast, P4 found the same setup isolating and unsettling, likening the stillness to lockdown loneliness and wishing for “more aromas or something” to enrich the experience.
@@
Others found the hood sharpened perception of detail. P5 noted, “It does help you focus on that, like this is what you’re eating.” P6 appreciated the privacy, “I didn’t really like…watching you eat bones… so it was nice to eat freely and not feel like people were watching me… it made the experience more intimate.” P3 described the most ambivalent reaction: feelings of “shame and guilt” as if engaging in something forbidden, mixed with curiosity and excitement. Crucially, she stressed that these tensions stemmed less from the food itself than from the performative act of concealment. For her, hiding under a napkin generated novelty, surprise, and playfulness alongside discomfort.
@@
These divergent responses show how ritualized interventions can shift social dynamics, provoking new avenues for reflection. The napkin ritual was not simply an immersive act: it reshaped sensory perception, interpersonal awareness, and ethical imagination.
By immersing participants in unfamiliar sensory worlds, the dinner-in-the-drama offered an embodied glimpse of futures that lies between plausible and preposterous (Wang et al., 2023a)—where biohybrid robots might be consumed to preserve or revive a dying tradition such as eating Ortolan Bunting. These experiences opened participants to explore other radical potentials: futures not only technologically novel but culturally and sensorially alien to current norms.
The combination of bodily intimacy and collective ritual created conditions for curiosity, discomfort, and shared reflection. These findings extend work in speculative design and experiential prototyping (Auger, 2013; Dunne and Raby, 2024; Michael, 2012; Jewitt et al., 2022), showing how eating as a medium uniquely foregrounds questions of ethics and materiality. Whereas many speculative methods rely on images, videos, or texts (e.g., (Alfrink et al., 2023; Simeone et al., 2022; Petsolari et al., 2024)), speculative dining demonstrates how tangibility and embodiment illuminate the politics of sensory experience in ways representational formats cannot.
Building on post-COVID design (Lupton and Willis, 2021) and sensory futures scholarship (Pink, 2023), we argue that both speculative and sensory methods can be combined (see also (Jewitt et al., 2022)) to illuminate wider social, technical, and political entanglements. Futures are not imagined in abstraction but experienced through bodies, memories, and collective rituals.
4.2.2. The Value of Cross-Media Theatricality
Participants reflected on how cross-media representations—including 3D-printed models, slideshow drawings, decorative birds, and live drones—shaped their perceptions of the boundary between “animal” and “machine”. Design cues strongly mediated engagement: P5 noted that the absence of eyes in the 3D prints “helped it remain an object”, while P4 preferred the hybrids “stay machine-like”. These choices reveal how the curation of the design fiction influenced participants’ speculations and critical reflections.
@@ 4.2.2. The Value of Cross-Media Theatricality
The hands-on, participatory character of the theatrical method was also generative. P6 described handling the models and watching the drone as a form of imaginative play: “I’m totally game for that… I can pretend this is a bird and not a drone… it led to a good moment where the drone kind of had a mind of its own.” P4 and P5 similarly emphasized that physically interacting with the models and drones stimulated richer speculation than written reports or static visuals, while group discussion further expanded perspectives. P2 extended this idea, framing food itself as a medium of speculation: “We were literally putting the meat from the presentation inside our bodies”—making imagined futures both tangible and memorable. As P4 summarized, such interactive experiences are “much more effective than one person sitting down with a laptop and typing out a scenario”.
Beyond embodied engagements with cross-media representations including material objects, participants highlighted how the theatrical setup deepened reflection. P1 likened the interplay of physical objects and live performance to theater, arguing that even malfunctions should be “embraced in character… never apologizing… everything is within the theater”. The staging, assigned roles, and curated narrative demanded role-play—and this was methodologically productive. P1 described inhabiting a dual position as “restaurant manager” and critical observer, performing enthusiasm while also questioning the technology. This was a blend of positionalities that stretched their thinking and imagination. P6 similarly slipped between his assigned persona—a 2050s Parisian filmmaker—and his own identity, an unsettling doubling that prompted deeper ethical reflection on biohybrid meat.
4.2.3. The Value of the Unexpected
@@
Participants highlighted how unexpected moments and exchanges within the dinner-in-the-drama shaped their experiences and provoked deeper reflections. P2 observed that the drone’s unplanned error during the demonstration “reminded us that machines don’t always do what we want them to do” evoking science fiction tropes where machines seem to possess minds and wills of their own. This moment underscored the limits of human control and illustrated how malfunction was not a limitation of the method but a catalyst, sparking reflection. Similarly, P5 described the drone’s erratic behaviour as a “quiet reminder that speculative technologies will inevitably go wrong”, emphasizing the importance of anticipating risks and strategies for response. For P3, however, the same incident was striking, describing the crash as “really beautiful”, suggesting that the machine’s uncontrolled movements imbued it with a sense of “soul”. Such unexpected moments can be seen to directly inform the generation of substantive themes unpacked in subsubsection 4.1.3.
@@
If drone malfunctions marked unexpected moments from the drama itself, equally significant were the unplanned contributions from participants as they role-played characters and interacted with one another. Group discussion was encouraged throughout and after the curated experience. P2 reflected that the act of eating together also provided a conducive social environment for exchanging viewpoints and speculation, creating imaginative setting that a written report or static interview could not achieve. This social environment, full of fast-paced lively exchanges, encroached on territories that would unlikely be given thought through individual contemplation.
Crucially, participants were invited to improvise with the narrative, props, and environment of the design fiction. This opened space for idiosyncratic references to literature, philosophy, and popular culture to surface. For example, P2 argued that biohybrid technologies push ethics into uncharted territory, where distinctions between complex machines and sentience begin to unravel (see subsubsection 4.1.3). To illustrate, he invoked Westworld and Ex Machina, noting how popular debates about machine ethics focus on humanoid robots while largely overlooking animal-like entities and biohybrids. Similarly, P4 referenced the Bicentennial Man to highlight how incremental human-like features shift perceptions of agency and moral worth.
@@
Together, these accounts frame unpredictability of the method not as disruption but as a speculative resource. While such moments cannot be fully designed or foreseen, the open, immersive, and performative nature of the dinner-in-the-drama created conditions where this productive serendipity could unfold—revealing how surprise, malfunction, and improvisation can deepen reflection on technological ethics and imagined futures.
The methodological reflections offered across this section position speculative dining as a site where future imagination and collective memory intersect. By staging encounters that are playful, unsettling, unpredictable, and socially charged, such practices push speculative design beyond abstract futuring: they invite publics to grapple with the ethical, affective, and cultural entanglements of technologies-in-the-making. Doing so the method activated cultural repertoires that bridged experiential design, philosophical inquiry, and everyday discourse. Rather than illustrating possible futures, speculative dining demonstrates how experiential prototyping can materially intervene in collective understanding, generating embodied memories and ethical awareness rarely achieved in conventional engagements. Our work advances design fiction from representation to embodiment, where multisensory encounters provoke unique reflections.
5. Conclusion and Future Work
This paper staged a tangible design fiction—a dinner-in-the-drama—where participants consumed “biohybrid flying robots” inspired by (but not in support of) the banned French delicacy of Ortolan Bunting. By moving speculation off the page and into embodied, multisensory experience, we invited publics to grapple with futures that are both provocative and unsettling.
Our findings highlight three interconnected terrains. First, technological feasibility and pathways (subsubsection 4.1.1) were imagined in ways that extended and brought together debates in AI, robotics and other areas of technological innovation. Participants drew on their embodied experiences of the tangible design fiction and wider knowledge to speculate about how this technology might be developed, adopted or rejected. They connected sensory qualities—such as taste and texture—with broader questions of production, distribution, and marketing. The nuanced accounts that surfaced reflect a set of opportunities and barriers often overlooked in conventional foresight. Second, cultural and ritual transformation (subsubsection 4.1.2) revealed how biohybrid food futures are never purely technical. Participants framed them through cultural politics, positioning biohybrids as means of preserving heritage practices, igniting new rituals, or provoking contestation. Their reflections revealed tensions between conservative impulses to adapt technologies to preserve existing traditions and progressive desires to craft novel culinary forms. Third, ethics and moral boundaries (subsubsection 4.1.3). Participants raised ethical questions at multiple levels: ontological (sentience, soul, suffering), perceptual (individuality, symbolism), and sociological (replicability, runaway risks). Resonating with posthumanist critiques, these reflections suggest that biohybrids may need to be understood as ethically significant others, especially as edible robotics move from lab to plate. Together, these terrains show how technological, cultural, and ethical reasoning intertwined when speculation was grounded in multisensory experience.
Beyond these thematic insights, our work demonstrates the methodological value of tangible, multisensory speculation. By integrating food, props, performance, and ritual, the dinner-in-the-drama fostered affective intensity, playful improvisation, and productive serendipity. These qualities encouraged participants to navigate uncomfortable questions—ones unlikely to emerge from textual scenarios alone. We argue that such methods enrich HCI’s repertoire for engaging publics with technologies-in-the-making, particularly when futures are materially intimate, culturally charged, and ethically fraught.
For HCI and HRI, our contributions are threefold: (i) a speculative artifact that stages robots as food, expanding imaginaries of human–robot interaction into culinary domains; (ii) empirical insights showing how non-technical experts negotiate feasibility, culture, and ethics when immersed in embodied speculation; and (iii) a methodological advance in tangible design fiction, extending speculative design toward experiential and multisensory provocation.
Looking forward, we encourage future work to explore speculative dining and related methods not only as entertainment but as critical tools for public engagement, ethical deliberation, and cross-disciplinary dialogue. As edible robotics and biohybrid systems move from fiction toward feasibility, the challenge for HCI is not simply to assess usability or efficiency but to grapple with how such technologies reconfigure cultural meaning, social relations, and ethical boundaries.

### v2

@@ 4.2. Tangible Multisensory Dining Experience as a Valuable Provocation
As shown above, our method and multisensory design surfaced a rich and productive range of perspectives and tensions. Participants consistently connected their technological speculations to deeper reflections on politics, religion, nationality, literary canons, symbolism, and ontological beliefs about the nature of life, consciousness, and suffering. In what follows, we explore three mechanisms through which these deep speculative exchanges emerged.
4.2.1. Immersive Engagement and Emotive Responses
@@
Participants felt that the immersive dining format—particularly the use of napkin hoods and live/mechanical elements—provoked strong sensory engagement and elicited intense emotional responses. Several highlighted the hooded moment as especially impactful, though their interpretations diverged. P2 described how covering heads with napkins transformed the room: conversation ceased and a collective silence emerged, “like putting a cover over a birdcage”. Rather than diminishing sociality, this paradoxically heightened his awareness of others: he became “hyper-aware of the other people eating around me”, sensing a subtle connection through shared quietude. By contrast, P4 found the same setup isolating and unsettling, likening the stillness to the loneliness of COVID-19 lockdowns.
@@
P5 found the hood sharpened perception of detail, including the textures and smells of food, “It does help you focus on that, like this is what you’re eating.” P6 appreciated the privacy, “I didn’t really like…watching you eat bones… so it was nice to eat freely and not feel like people were watching me… it made the experience more intimate and helped me really focus on the act of eating.” P3 described the most ambivalent reaction: feelings of “shame and guilt” as if engaging in something forbidden, mixed with curiosity and excitement. Crucially, she stressed that these tensions stemmed less from the food itself than from the performative act of concealment. For her, hiding under a napkin generated novelty, surprise, and playfulness alongside discomfort.
@@
Overall, these divergent responses show how ritualized interventions can shift social dynamics and reshape sensory perception, interpersonal awareness, and ethical imagination, provoking new avenues for reflection.
4.2.2. Multisensory Cross-Media Theatrical Design
Participants reported that cross-media representations, including slideshow drawings, decorative birds, the live drone, the 3D-printed model, and the edible roasted birds shaped their perception of the boundary between “animal” and “machine”.
Design cues strongly mediated anthropomorphism and engagement: P5 observed that the absence of eyes in the model helped it remain object-like: “The moment you have eyes, I think, makes it a person… the absence of that helped.” P1 likened the 3D print to a “butchered” version of the bird-drone, highlighting contrasts between living and mechanical forms. P4 emphasized the robot-like qualities of the hybrid, noting its mechanical parts and the design logic of starting with a robot and adding biological material; he noted how gradual addition of human features can shift perception but preferred the hybrid to remain clearly machine-like, reinforcing its role as a meat-production device.
P2 noted that the cross-media layering representations suggested possibilities for future iterations, imagining a synthesis of form and function: “I could imagine… with more resources… to have the drone technology that makes it function join with that beautiful 3D printed resin model.” He further suggested that adding feathery textures on the drone model could enhance perceived vitality, noting that even the small decorative birds (see Figure 9b) on the table contributed a stronger sense of life and observation. Participants also noted that seeing slideshow representations (e.g., Figure 4b) of how animals are raised, prepared, and consumed clarified both practical and ethical aspects.
@@ 4.2.2. Multisensory Cross-Media Theatrical Design
The multisensory hands-on, participatory character of the theatrical method was also generative. P6 described handling the models and watching the drone as a form of imaginative play: “I’m totally game for that… I can pretend this is a bird and not a drone… it led to a good moment where the drone kind of had a mind of its own.” P4 and P5 similarly emphasized that physically interacting with the models and drones stimulated richer imagination and reflection than written reports or static visuals. P2 extended this idea, framing the dining itself as a medium of engagement: “We were literally putting the meat from the presentation inside our bodies”—making imagined futures both tangible and memorable. P3 added that “It felt like we were acting in a sci-fi movie.” As P4 summarized, such interactive experiences are “much more effective”, while P2 went further, positioning play as the key mechanism:
“This is a wonderful way to help people engage with a speculative future… there was a situation that allowed us all to dwell and process in it.”
4.2.3. The Serendipity of Unexpectedness and Improvisation
All participants noted that an unexpected drone malfunction profoundly shaped their experience of the drama and provoked deeper reflections. During the demonstration, insufficient ambient lighting caused the drone’s vision system to fail, sending it off course and into the corner of the room. This brief “out-of-control” moment became a focal point of the experience; Figure 4c captures participants’ reactions.
“I especially like the part where it tried to escape… well, it made it more real… it’s trying to get away.” (P4) “That made it much stronger for the guests and much more thought-provoking because then this thing had its own free will.” (P1) “It reminded us that machines sometimes malfunction and that humans aren’t always perfect at making things go how we want with our technology.” (P2)
@@ 4.2.3. The Serendipity of Unexpectedness and Improvisation
Similarly, P5 described the erratic drone behaviour as a “reminder that speculative technologies will inevitably go wrong”, emphasizing the importance of anticipating risks and strategies for response. For P3, the same incident was striking, describing the crash as “really beautiful”, suggesting that the uncontrolled movements imbued the machine with “a sense of soul”. Such unexpected moments can be seen to directly inform the generation of substantive themes unpacked in 4.1.2.
@@
While drone malfunctions introduced unexpected moments into the drama, equally significant were participants’ unplanned contributions as they role-played and interacted with one another. Group discussion was encouraged throughout and after the curated experience. P2 reflected that the act of eating together also provided a conducive social environment for exchanging viewpoints and speculation, creating an imaginative setting that an individual alone could not achieve. This social environment, full of fast-paced lively exchanges, encroached on territories that would unlikely be given thought through individual contemplation.
Participants were invited to improvise with the narrative, props, and environment of the design fiction. This opened space for idiosyncratic references to literature, philosophy, and popular culture to surface. For example, P2 connected the concept of creating biohybrid robot meat to Shakespeare’s line, “the green-eyed monster which doth mock the meat it feeds on”, commenting that the dinner theater was effectively “mocking meat”. He used the pun to suggest that the edible robot enactment both imitated natural meat and satirized the problems of contemporary meat production. This dual meaning highlights a paradox: alternatives designed to be ethically superior may still mimic the very practices they aim to replace.
@@
Together, these accounts frame unexpectedness and improvisation not as disruption but as a speculative and provocative resource. While such moments cannot be fully designed or foreseen, the open, immersive, and performative nature of the dinner-in-the-drama created conditions where this productive serendipity could unfold—revealing how surprise, malfunction, and improvisation can deepen reflection on technological ethics and imagined futures.
5. Discussion
5.1. Ambiguity in Lay Imaginaries of Biohybrid Foods
Participants’ future imaginaries of biohybrid foods were dynamic, oscillating, and sometimes contradictory. They speculated about timelines, market adoption, taste, sustainability, and social impacts, revealing a complex landscape of hopes, anxieties, and ethical considerations. Rather than treating these oscillations as misunderstandings or gaps in knowledge, we interpret them as reflections of the inherent ambiguity of biohybrid foods. Prior speculative proposals around robot meat and edible robotic systems suggest that this ambiguity is not unique to our study but characteristic of broader imaginaries of hybridized food–machine entities (Wang et al., 2023a; Floreano et al., 2024; Zhang et al., 2025). Drawing on Gaver’s concept of ambiguity as a resource (Gaver et al., 2003) and Barad’s notion of indeterminacy (Barad, 2007), we conceptualize biohybrid foods as inherently ambiguous technologies: entities that actively blur boundaries between biological and mechanical, alive and designed, edible and symbolic.
The speculative dining format amplified this ambiguity by foregrounding sensory and embodied provocation. Taste, texture, sound, and ingestion rituals encouraged participants to model critical scenarios of adoption, asking how biohybrids might enter everyday life, be regulated, or be personalized. This embodied engagement enabled participants to explore future imaginaries and ritual transformation, imagining how culinary practices, cultural rituals, and ethical norms might shift in response to hybridized foods.
At the same time, several forms of ambiguity emerged through unscripted and unexpected moments during the dining encounters. These moments were not designed as fixed narrative prompts; they arose contingently through sensory surprise, improvisation, or unplanned technical errors. In line with arguments that ambiguity invites multiple interpretations and sustained personal engagement (Gaver et al., 2003), these unpredictable encounters prevented any single interpretive frame from stabilizing. Such “productive indeterminacy” is not a deficit of knowledge to be resolved, but an ongoing material–discursive process in which boundaries come to matter through practice (Barad, 2007). In this sense, ambiguity was not merely present but intensified through emergent, unanticipated interactions with speculative artifacts.
The oscillating and situated nature of these imaginaries shows that ambiguity is not a barrier to understanding; it is a productive epistemic and design space that fosters ethical imagination, collective sense-making, and culturally grounded speculation. Through this lens, we contribute to work in HCI, HRI, and HFI by identifying ambiguity itself as an empirical pattern in lay imaginaries of emerging food technologies. Rather than treating oscillation as noise or confusion, we show it to be a structuring feature of how people make sense of biohybrid futures through sensory, ethical, and affective reasoning. This subsection therefore offers a conceptual shift: indeterminacy is not simply a theoretical property of hybrid organisms, but a situated cognitive and social practice enacted through engagement with ambiguous artifacts (Gaver et al., 2003; Barad, 2007).
5.2. Edibility as a Negotiated Ethical Boundary
Participants’ encounters highlighted edibility as a central and negotiated ethical boundary. Rather than a fixed property, edibility emerged as a boundary object connecting biology, technology, culture, and personal experience (Star and Griesemer, 1989; Probyn, 2016; Paxson, 2013; Bingham, 2006). Questions surfaced regarding conditions under which consuming hybrid organisms might be acceptable, and what forms of responsibility are owed to semi-living entities. These concerns echo ongoing debates in biohybrid robotics ethics, where researchers argue that boundaries between organism, technology, and moral obligation are negotiated rather than given (Mestre et al., 2024; Kirksey, 2021). As others note, imaginaries of “living machines” often draw from cultural narratives of control over life (Vaage, 2020), and here such narratives are refracted through the intimate act of ingestion.
Ethical reflection was embodied and relational. Taste, texture, and movement shaped judgments that oscillated between attraction and aversion, curiosity and caution. Participants invoked frames from synthetic biology, AI ethics, and religious belief (Hyun et al., 2016; for Stem Cell Research, 2021; Webster-Wood et al., 2022) to reason about sentience, individuality, and replicability. In this way, biohybrid foods prompted situated renegotiation of “edible versus inedible”, making moral and cultural boundaries explicit and mutable. Viewed through posthumanist perspectives (Braidotti, 2013; Haraway, 2008), these reflections foreground human–nonhuman entanglements, where ethical reasoning arises through encounter with semi-living, partially autonomous entities rather than abstract principle.
Speculative dining functioned as a method for ethical boundary-work, enabling participants to experiment with norms, values, and social practices around hybrid foods. Instead of resolving whether biohybrids should be eaten, the dining encounters surfaced how technological, sensory, cultural, and ethical considerations co-produce what counts as food, and under what conditions. For HCI and HFI, this positions edibility not as an inherent attribute of objects, but as a relational and negotiated property continually shaped through sensory interaction and shared deliberation.
5.3. The Cultural and Affective Politics of Synthetic Eating
Participants’ reactions to the speculative dishes demonstrated that synthetic eating is not merely an ethical or sensory practice, but deeply embedded in cultural values and affective frameworks. Their interpretations invoked memories of family traditions, religious norms, regional culinary identities, and deeply held taboos—showing that biohybrid foods enter cultural fields already structured by history, identity, and emotional attachments. As previous work in food politics argues, eating practices are inseparable from collective identity, moral order, and social meaning (Counihan, 2013; Fischler, 1988), and our results confirm that participants mobilized such frameworks to make sense of hybridized foods.
Throughout the dining sequence, individuals continuously mapped unfamiliar specimens onto familiar cultural categories, echoing anthropological accounts of how novel or “exotic” foods are domesticated through analogy and metaphor (Sutton, 2006; Fischler, 1988). These mappings were not superficial analogies: they materially shaped participants’ willingness to engage, emotional responses, and their narratives about potential futures. For example, when a hybrid specimen was associated with an existing ritual food, some participants considered it acceptable for ceremonial consumption; when another evoked tabooed species, unease or refusal followed. This pattern aligns with research in Human–Food Interaction showing that sensory and cultural cues (e.g. smell, texture, appearance) significantly influence the perceived acceptability of unfamiliar food–technology hybrids (Obrist et al., 2014b, 2016).
Beyond cognitive mapping, the experience triggered a spectrum of affective responses—curiosity at novelty, discomfort at blurred organism–machine boundaries, protectiveness toward malfunctioning drones or “fragile” tissue samples, and even excitement about future possibilities. These emotional reactions played a structuring role: discomfort often led to rejection, curiosity prompted further engagement, and care sometimes reframed hybrids as quasi-relational entities deserving ethical consideration. This resonates with scholarship that positions emotion and affect not as background reactions but as central in stabilizing—or destabilizing—categories of edibility, purity, and moral acceptability (Probyn, 2016; Fischler, 1988). In our context, affective engagement was facilitated through material and interactive design—for instance, by invoking uncertainty, fragility, or familiarity through multisensory artifacts, akin to how edible-robotic or microbial food-technology work uses sensory design to elicit intimacy, care, or repulsion (Floreano et al., 2024; Chen et al., 2021).
Participants also frequently invoked broader social norms and anticipated public reactions—imagining how others might respond, speculating on collective acceptance or resistance, and foreseeing that hybrid foods could trigger new rituals or social divisions. This echoes classic anthropological theory that food practices act as cultural boundary markers (Douglas, 1972). In our study, synthetic eating triggered three types of boundary-work: (i) reinforcing existing cultural distinctions (e.g. natural vs. unnatural, sacred vs. profane), (ii) reworking traditions by imagining hybrid meals that combine conventional and novel elements, and (iii) projecting future cultural orders in which hybrid foods serve as markers of ecological adaptation, social status, or technological modernity. Similar projections appear in speculative design and food-futures literature, where imagined consumption practices become sites of social imaginaries and value negotiation (Dunne and Raby, 2024; Chopra et al., 2022).
Together, these findings highlight that adopting biohybrid foods cannot be understood solely as a question of technical feasibility or individual taste—it is fundamentally a process negotiated through cultural frameworks, affective attachments, and imagined social futures. For HCI, HFI, and HRI research concerned with emerging food technologies, these results underscore the importance of attending not only to design and usability, but also to the cultural infrastructures and affective economies through which hybrid foods may acquire meaning and legitimacy.
5.4. Tangible Design Fiction as Ethical Sandbox and Public Co-Theorization
Speculative dining functioned as a tangible form of design fiction in which ethical and ontological questions were enacted through ritualized, multisensory practice rather than merely discussed in the abstract. Hooding, cross-media staging, and direct engagement with hybrid food objects generated designed disruptions—moments of curiosity, discomfort, and collective attentiveness that foregrounded shifting relations of visibility, agency, and responsibility (Auger, 2013; Michael, 2012; Jewitt et al., 2022). These interventions did not simply represent imagined futures; they staged conditions for experimenting with ethical boundaries in situ. This extends prior work on speculative design showing that embodied, non-representational encounters surface ethical tensions that remain less visible in textual or visual design fiction (Dunne and Raby, 2024; Alfrink et al., 2023; Simeone et al., 2022; Petsolari et al., 2024). Within HFI, our approach builds on edible speculative interfaces and multisensory provocations (Grimes and Harper, 2008; Kato et al., 2024; Song et al., 2024; Deng et al., 2023b, 2024), suggesting that tangible encounters can serve as provisional “test sites” where hybrid food practices are rehearsed before they exist at scale.
Crucially, these encounters operated as ethical sandboxing rather than controlled evaluation. Participants developed and revised interpretive frameworks by drawing on market logics, ritual norms, AI ethics, posthumanist reasoning, and cultural taboos to model possible food futures (Jewitt et al., 2022; Lupton and Willis, 2021). In this mode, they were not merely reacting as informants; they were experimenting with ethical positions, imagining regulatory regimes, and anticipating cultural adaptations. Prior work on speculative participation and anticipatory governance suggests that publics can actively surface normative assumptions and propose alternative futures (Chopra et al., 2022; Ringfort-Felner et al., 2023; Jasanoff and Kim, 2015; Forlano and Mathew, 2014). Our findings resonate with this work: the micro-negotiations around responsibility (“who is accountable if a semi-living food system malfunctions?”), ritual (“how would hybrid foods be blessed, shared, refused?”), and agency (“does a semi-autonomous edible robot have a right not to be consumed?”) show publics performing ethical reasoning normally reserved for expert domains.
Taken together, speculative dining enacted a form of collective, embodied sandboxing in which participants tested moral intuitions, rehearsed social norms, and learned through sensory and relational trial-and-error. Rather than producing consensus, these interactions generated provisional, situational insights grounded in lived experience. This suggests a methodological implication for HCI, HRI, and HFI: when publics are engaged through sensory, material, and shared encounters, they co-produce anticipatory knowledge about emerging food technologies in ways that are exploratory, reversible, and ethically consequential. In this sense, the sandbox is not metaphorical but material—structured through the textures, movements, and rituals of speculative eating—and it enables publics to co-theorize futures in ways that are iterative, uncertain, and open-ended.
6. Limitations and Future Work
Our participant group was small and intentionally composed of individuals from creative and cultural sectors. This supported depth of reflection but limits the transferability of the findings. To address this, future work should engage broader publics—including culinary practitioners, technologists, ethicists, and general consumers—to examine how diverse communities interpret speculative biohybrid food systems and how speculative encounters vary across social worlds.
The scenario drew heavily on the culturally specific symbolism of the Ortolan ritual within a fictional 2052 Parisian setting. This framing likely influenced the meanings participants attributed to ritual, taboo, and hybrid organisms. Future studies could adapt the dinner-in-the-drama to different cultural contexts, exploring how food traditions, religious norms, and local symbolic systems shape public engagement with synthetic and hybrid forms of eating.
Although our findings highlight how multisensory tangible design fiction can support ethical and cultural reasoning, the study did not compare this method against alternative speculative modalities. To determine what is unique about embodied, sensory engagement, future research should systematically compare dinner-in-the-drama experiences with text-based scenarios, video design fictions, VR environments, or non-sensory roleplay to assess how different modalities elicit different forms of reflection.
Our speculative artifacts were supported by Wizard-of-Oz control and culinary proxies rather than functioning biohybrid systems. These choices ensured feasibility and participant safety but may have shaped perceptions of agency, materiality, and realism. Future work could explore higher-fidelity or interactive prototypes—such as soft robotics, edible robotics, bioinspired robotics (e.g., a flapping-wing drone (Wang et al., 2025a)), biohybrid robotics, or tissue-mimicking constructs—to investigate how increasing realism affects ethical reasoning and sensory interpretation.
7. Conclusion
This paper presented a tangible design fiction—a multisensory dinner-in-the-drama—that invited participants to confront the imagined consumption of edible biohybrid flying robots. By shifting speculation from narrative description to embodied encounter, the work revealed how technological, cultural, and ethical reasoning becomes intertwined when futures are experienced through taste, touch, sound, and ritual rather than imagined at a distance. Participants used the sensory and symbolic affordances of the scenario to speculate on technological pathways, reinterpret culinary traditions, and negotiate moral boundaries around hybrid organisms, surfacing perspectives often overlooked in conventional foresight. This work expands HCI’s repertoire for engaging publics with emergent biotechnologies, especially those poised to reshape intimate, everyday practices such as eating. As edible robotics and biohybrid systems might move closer to feasibility, the challenge for HCI is not only to evaluate their functionality but to understand how they will transform cultural meaning, social relations, and ethical worlds.